Sun Microsystems is a leading supplier of computer-related products, including servers, workstations, storage devices, and network switches. In 2009, Sun Microsystems was acquired by Oracle Corporation. In the letter to stockholders as part of the 2001 annual report, President and CEO Scott G. McNealy offered the following remarks: Fiscal 2001 was clearly a mixed bag for Sun, the industry, and the economy as a whole. Still, we finished with revenue growth of 16 percent—and that's significant. We believe it's a good indication that Sun continued to pull away from the pack and gain market share. For that, we owe a debt of gratitude to our employees worldwide, who aggressively brought costs down—even as they continued to bring exciting new products to market. The statement would not appear to be telling you enough. For example, McNealy says the year was a mixed bag with revenue growth of 16 percent. But what about earnings? You can delve further by examining the income statement in Exhibit 4. Also, for additional analysis of other factors, consolidated balance sheet(s) are presented in Exhibit 5.

2001 Dollars 2000 Dollars 1999 Dollars 1998 Dollars
Net revenues $ 18,325 $ 15,715 $ 11,803 $ 9,889
Costs and expenses:

     Cost of sales

$ 10,045

$ 7,548

$ 5,671

$ 4,650

    Research and development

2,012

1,627

1,273

1,033

    Selling, general and administrative

4,540

4,061

3,198

2,835

    Goodwill amortization

264

63

17

0.6

    In-process research and development

81 12 117 179
Total costs and expenses $ 16,942 $ 13,311 $ 10,276 $ 8,697.6
Operating Income $ 1,383 $ 2,404 $ 1,527 $ 1,191.4
Gain (loss) on strategic investments $ -87 $ 214 - -
Interest income, net $ 360 $ 167 $ 86 $ 47
Litigation settlement - - - -
Income before taxes $ 1,656 $ 2,785 $ 1,613 $ 1,238.4
Provision for income taxes $ 817.24 $ 922.7 $ 664.41 $ 539.24
Cumulative effect of change in accounting principle, net $ -52 - - -
Net income $ 890.76 $ 1,862.3 $ 948.59 $ 699.16
Net income per common share-diluted $ 0.26 $ 0.55 $ 0.29 $ 0.22
Shares used in the calculation of net income per common share-diluted 3,426 3,386 3,271 3,178

Exhibit 2

Assets 2001 2000
Current assets:
     Cash and cash equivalents $ 1,478 $ 1,858
     Short-term investments 387 621
     Accounts receivable, net allowances of $410 in 2001 and $534 in 2000 2,954 2,683
     Inventories 1,045 559
     Deferred tax assets 1,089 679
     Prepaids and other current assets 973 479
          Total current assets 7,926 6,879
Property, plant and equipment, net 2,696 2,093
Long-term investments 4,675 4,486
Goodwill, net of accumulated amortization of $349 in 2001 and $88 in 2000 2,037 163
Other assets, net 831 525
18,165 14,146
Liabilities and Stockholders' Equity
Current liabilities:
     Short-term borrowings 4 9
     Accounts payable 1,042 923
     Accrued payroll-related liabilities 490 750
     Accrued liabilities and other 1,374 1,150
     Deferred revenues and customer deposits 1,821 1,294
     Warranty reserve 317 211
     Income taxes payable 93 203
          Total current liabilities 5,141 4,540
Deferred income taxes 750 572
Long-term debt and other obligations 1,705 1,717
Total debt 7,596 6,829
Commitments and contingencies
Stockholders' equity:
     Preferred stock, $0.001 par value, 10 shares authorized (1 sahre which has been designated as Series A Preferred participating stock): no shares issued and outstanding - -
Common stock and additional paid-in-capital, $0.00067 par value, 7,200 shares authorized; issued: 3,536 shares in 2001 and 301 shares in 2000 6,235 2,732
Treasury stock, at cost: 288 shares in 2001 and 301 shares in 2000 -2,433 -1,442
Deferred equity compensation -75 -17
Retained earnings 6,872 5,966
Accumulated other comprehensive income (loss) -30 78
     Total stockholders' equity 10,569 7,317
18,165 14,146

Part A

Referring to Exhibit 1, compute the annual percentage change in net income per common share-diluted (second numerical line from the bottom) for 1998–1999, 1999–2000, and 2000–2001.

Rate of change, 1998 to 1999: _________

Rate of change, 1999 to 2000: _________

Rate of change, 2000 to 2001: _________

Part B

Also in Exhibit 1, compute net income/net revenue (sales) for each of the four years. Begin with 1998.

1998 Profit Margin: __________

1999 Profit Margin: __________

2000 Profit Margin: __________

2001 Profit Margin: __________

Part C

Compute return on stockholders’ equity for 2000 and 2001 using data from Exhibits 1 and 2.

2000 Return on Stockholders' Equity: _______

2001 Return on Stockholders' Equity: _______

Part D

Analyze your results to Question 2 (Part B above) more completely by computing ratios 1, 2a, 2b, and 3b (these numbers correspond with the numbered ratios listing in Chapter 3 of our textbook) for 2000 and 2001. Actually, the answer to ratio 1 can be found as part of the answer to question 2 (Part B), but it is helpful to look at it again.

Ratio 2000 2001
1
2a
2b

Part E

The average stock prices for each of the four years shown in Exhibit 1 were as follows:

1998          11¼

1999          16¾

2000          28½

2001          9½

Compute the price/earnings (P/E) ratio for each year. That is, take the stock price shown above and divide by net income per common stock-dilution from Exhibit 1.

1998 P/E Ratio: ________

1999 P/E Ratio: ________

2000 P/E Ratio: ________

2001 P/E Ratio: ________
Answer #1

PART A

Rate of change, 1998 to 1999: 0.29-0.22/0.22=31.82%

Rate of change, 1999 to 2000: 0.55-0.29/0.29= 89.66%

Rate of change, 2000 to 2001: 0.26-0.55/0.55= -52.73%

PART B

1998 Profit Margin: 699.16/9889=7.07%

1999 Profit Margin: 948.59/11803= 8.04%

2000 Profit Margin: 1862.3/15715= 11.85%

2001 Profit Margin: 890.76/18325= 4.86%

PART C

2000 Return on Stockholders' Equity: Net income/ Stockholders' equity= 1862.3/7317= 25.45%

2001 Return on Stockholders' Equity: 890.76/10569=8.43%

PART E

1998 P/E Ratio: 11.25/0.22=51.14

1999 P/E Ratio: 16.75/0.29=57.76

2000 P/E Ratio: 28.5/0.55=51.82

2001 P/E Ratio: 9.5/0.26=36.54